# Performance and Generalizability Impacts of Incorporating Geolocation into Deep Learning for Dynamic PM2.5 Estimation

### v1

@@ 3.2 Model Architecture
    where $\odot$ denotes Hadamard product. This operation allows the model to learn the interactions between temporal patterns and geographic context. The fused embedding $e_{\text{fused}}$ is then passed to a final fully connected layer that regresses the surface-level PM2.5 concentrations, allowing the model to rely on both the time series observations as well as pre-learned embeddings for the given location.
3.3 Model Training
@@ 4 Results
We assess model performance and generalizability using three main evaluation settings, followed by ablation and qualitative analyses. In all experiments, we report the coefficient of determination ($R^{2}$) and root mean square error (RMSE) as performance metrics for average of five training runs. Please note that the results within each setting can be compared, not across settings, to ensure the number of training samples, spatial distributions, and all other configurations remain similar.
@@ 4.4 Qualitative Analysis of Spatial Patterns
While the summary performance metrics provided in the previous sections shed light on the overall estimation performance of the models, we can visualize the model output over space. However, it is worth remembering that there is no ’ground truth’ available for the entire space, i.e., there is no satellite can observe the surface-level PM2.5 concentrations at all locations against which the model output can be compared. Figure fig. 6 provides a comparison of the maps generated using two different models, one without any geolocation feature (a), and (b) with pretrained GeoCLIP features, which performed best in the OOR and WR evaluation scenarios.
@@ 5.1 Estimation Performance Comparison
We built on a previously validated Bi-LSTM with Attention model that achieves state-of-the-art performance in PM2.5 estimation, including during high-concentration events [27]. Compared to the Di et al. [30] dataset from 2005 to 2016, our model improves RMSE from 2.73 to 2.63 (--3.7%) for all days, and 19.01 to 15.44 (--18.8%) for high concentration days ($>$35 $\mu$g/m3). Compared to Wei et al. [31] dataset which covers 2017 to 2021, for the same period, our base Bi-LSTM model improves the RMSE from 4.70 to 2.73 (--41.9%) for all days, and 26.45 to 20.78 (--21.4%) for high concentration days ($>$35 $\mu$g/m3). Our Bi-LSTM with Attention explains 10% more variance than our own Random Forest strong baseline, and 4% more than a regular LSTM; RMSE is 3.59 (vs RF’s 4.00, --10.3%, and LSTM’s 3.81, --5.8%). More details and comparisons of the underlying air pollution estimation model can be found in [27].
@@ 5.2 On the Impact of Coordinate Transformation Methods
Another question might be whether more expressive strategies like Fourier feature expansion [49]—would improve model performance, and more importantly, geographic generalizability. Let us consider Fourier feature mappings, which encode scalar inputs using a series of trigonometric functions at increasing frequencies:
$\gamma(s)=\left[\sin(2^{0}\pi s),\cos(2^{0}\pi s),\sin(2^{1}\pi s),\cos(2^{1}\pi s),\ldots,\sin(2^{k-1}\pi s),\cos(2^{k-1}\pi s)\right]$
for some scalar $s$ (e.g., latitude or longitude) and user-defined number of frequencies $k$. These higher-frequency components enable neural networks to learn high-frequency variations (i.e., higher spatial resolutions) in the input space by making the mapping $f(\gamma(s))$ more expressive, compared to low-frequency encodings like:
$[\sin(\varphi),\cos(\varphi),\sin(\lambda),\cos(\lambda)]$
@@ 6 Conclusion
While the original intention of our work was to quantify the impacts of incorporating geolocation information into deep learning for a the estimation of a target with day-to-day changes with multi-variate observations, our results also highlight the value of location encoders (despite and in addition to the multi-variate domain observations), and the value of an emerging research body to distill information about the Earth in latent embeddings for seamless incorporation into deep learning. This area of research is relatively young, and so far, has been focused on fusing one mode of data with location representations. To distill even more information into these models, future research can move towards multi-modal and multi-sensor enhancements that better capture information about a place into these latent embeddings. As evidenced by our dynamic application that requires daily observation of aerosol content in the atmosphere, adding a temporal aspect to location encoders can also be invaluable in applications on dynamic phenomena.
@@
All data uses are openly available at https://doi.org/10.6084/m9.figshare.29139464.v2

### v2

@@ 2.3.2 Model Architecture
    where $\odot$ denotes Hadamard product. This operation allows the model to learn the interactions between temporal patterns and geographic context. The fused embedding $e_{\text{fused}}$ is then passed to a final fully connected layer that regresses the surface-level PM2.5 concentrations, allowing the model to rely on both the time-series observations as well as pre-learned embeddings for the given location.
2.3.3 Model Training
@@ 3 Results
We assess model performance and generalizability using three main evaluation settings, followed by ablation and qualitative analyses. In all experiments, we report the coefficient of determination ($R^{2}$), root mean square error (RMSE), and mean bias error (MBE) as performance metrics for average of five training runs. Complete evaluation tables are provided in Appendix Table 2-5. We provide key metrics in the following Sections. Please note that the results within each setting can be compared, not across settings, to ensure the number of training samples, spatial distributions, and all other configurations remain similar.
@@ 3.1 Within-Region (WR) Evaluation
Another noteworthy observation is that, under the random test scenario, GeoCLIP with Hadamard fusion yields the smallest bias (lowest $|\mathrm{MBE}|$), followed by the model without lat/lon (Table 2). Models using direct coordinate features (raw lat/lon or their sinusoidal transforms) exhibit higher bias, although none show practically significant systematic over- or under-estimation. Under the spatial split, the model without lat/lon achieves the smallest bias, while the model with GeoCLIP embeddings fusion display a slightly larger average bias, but with notably smaller standard deviations across splits, indicating more stable performance across spatial partitions (Table 3).
@@ 3.2 Out-of-Region (OoR) Evaluation
Across both checkerboard partitions, especially for $\delta{=}16^{\circ}$, GeoCLIP with Hadamard fusion maintains the smallest bias with similar values between the two held-out folds (partitions), whereas direct coordinate features lead to larger and more asymmetric biases (Tables 45). This further shows that pretrained location embeddings enhance spatial generalization, consistent with other findings.
To further assess the model’s geographic generalizability under a more extreme spatial configuration, we conducted an additional experiment using a checkerboard partition with $\delta=30^{\circ}$. This partitioning effectively divides CONUS into western and eastern halves. However, this configuration presents two key caveats. First, incorporating raw or sinusoidal geographic coordinates (e.g., latitude and longitude) in such a setting is inappropriate. Specifically, the longitude values in the test region become entirely out-of-distribution (OoD) relative to the training region, undermining the utility of direct coordinate-based features (our checkerboard evaluations earlier was partially motivated to prevent this from occurring). As a result, we do not train models with naive or sinusoidal lat/lon inputs in this evaluation. Second, PM2.5 dynamics exhibit well-known and stark differences between eastern and western CONUS, most notably, due to differing emissions profiles and the prevalence of wildfire smoke in the West. Therefore, in addition to geographic coordinates being out of distribution, the target value follows different dynamics too. While understanding these caveats, we can still compare models with no geolocation features against the ones enhanced with fused GeoCLIP embeddings.
[…]
@@ 3.4 Qualitative Analysis of Spatial Patterns
While the summary performance metrics provided in the previous sections shed light on the overall estimation performance of the models, we can visualize the model output over space. However, it is worth remembering that there is no ’ground truth’ available for the entire space, i.e., there is no satellite can observe the surface-level PM2.5 concentrations at all locations against which the model output can be compared. Figure 6 provides a comparison of the maps generated using two different models, one without any geolocation feature (a), and (b) with pretrained GeoCLIP features, which performed best in the OOR and WR evaluation scenarios.
@@
3.5 Evaluation Under High-Concentration Conditions: The 2021 Dixie Fire
The 2021 Dixie Fire was one of the largest and most destructive wildfires in California history, igniting on July 13, 2021 in the northern Sierra Nevada, near the Cresta Dam in Butte County. It scorched approximately 963,000 acres across five counties (Butte, Plumas, Tehama, Lassen, and Shasta). The fire destroyed nearly 1,300 structures and generated sustained high levels of PM2.5 across large areas of the western United States. Figure 8 compares the daily PM2.5 estimations of the baseline model (without geolocation features) and the GeoCLIP-enhanced model over four dates spanning peak fire activity. Notably, the GeoCLIP-enhanced model produces larger, contiguous plumes with higher PM2.5 concentrations over Northern California, better aligning with fire-affected areas. In contrast, the baseline model exhibits more spatial fragmentation and muted intensity, underestimating the scale and magnitude of pollution peaks. Interestingly, the GeoCLIP-enhanced model also captures elevated PM2.5 concentrations over northern Minnesota after July 21, coinciding with the documented regional smoke event caused by wildfires in both western North America and Canada [57, 58] at the time. These transported plumes led to air-quality alerts across the Upper Midwest, a pattern that the baseline model largely fails to reproduce.
We acknowledge the persistence of potential artifacts, for instance, along NW to SE direction for the GeoCLIP-enhanced model, or in the case of the fourth row of Figure 8, the relatively high values (light yellow) in disconnected patches surrounding the fire centers. In comparison, baseline model predictions look smoother across space, however, at the potential cost of underestimation at some high concentration areas.
Figure 8: Estimated PM2.5 during the 2021 Dixie Fire (Northern California) using baseline (no geographic features) and GeoCLIP-enhanced models. Each row shows a different date during peak wildfire activity. Columns (a) and (b) depict the baseline model output, while columns (c) and (d) show the GeoCLIP-enhanced model output at CONUS and regional scales. The GeoCLIP-enhanced model produces stronger plume intensity and spatial coherence over fire-affected regions and also captures elevated PM2.5 concentrations over northern Minnesota on July 21, consistent with documented long-range smoke transport from concurrent western U.S. and Canadian wildfires.
[…]
Figure 10: Residuals of PM2.5 estimations (model estimate minus observed) during the 2021 Dixie Fire, with NOAA HMS smoke plumes overlaid. Baseline model which does not use geographic features (left) shows systematic underprediction (blue) in fire regions. GeoCLIP-enhanced model (middle) reduces negative residuals, and aligns better with the smoke plumes. Rightmost column shows HMS smoke density for visual reference.
Figure 10 further visualizes the spatial residuals for the same dates alongside NOAA HMS smoke density polygons. The GeoCLIP-enhanced model reduces systematic negative bias in smoke-affected regions, as seen in the transition from large blue residuals (underprediction) in the baseline model to more neutral residuals. This indicates an improved capacity to capture elevated pollution levels during dynamic wildfire events. These observed patterns are consistent with the previous summary metrics and single-day maps: the GeoCLIP-fused model not only improves generalization across space, but also enhances larger value estimates under extreme air quality conditions. By encoding location semantics via pretrained embeddings, the model is better positioned to estimate higher concentrations in regions lacking dense observations but exhibiting complex pollution dynamics—such as those during wildfire plumes.
Despite improved performance, the GeoCLIP-enhanced model still underpredicts PM2.5 in smoke-affected areas. This residual bias may stem from limited downstream training data in extreme pollution conditions, the static nature of GeoCLIP embeddings (which cannot reflect evolving fire behavior or atmospheric transport), and uneven upstream pretraining coverage. Fine-scale variability due to complex terrain or plume lofting (e.g., over the Sierra Nevada) may also be poorly captured by models relying on coarser inputs. These results suggest that while pretrained location embeddings improve spatiotemporal estimates, further gains may require integrating further contextual features into location encoders, as well as more strategies for enhancing extreme value estimations in downstream training.
@@ 4.1 Estimation Performance Comparison
[…]
@@ 4.2 On the Impact of Coordinate Transformation Methods
Another question might be whether more expressive strategies like Fourier feature expansion [59]—would improve model performance, and more importantly, geographic generalizability.
The higher-frequency components of Fourier Expansion enable neural networks to learn high-frequency variations (i.e., higher spatial resolutions) in the input space by making the mapping more expressive, capturing higher resolution variations.
@@ 5 Conclusion
While the original intention of our work was to quantify the impacts of incorporating geolocation information into deep learning for the estimation of a target with day-to-day changes with multi-variate observations, our results also highlight the value of location encoders (despite and in addition to the multi-variate domain observations), and the value of an emerging research body to distill information about the Earth in latent embeddings for seamless incorporation into deep learning. This area of research is relatively young, and so far, has been focused on fusing one mode of data with location representations. To distill even more information into these models, future research can move towards multi-modal and multi-sensor enhancements that better capture information about a place into these latent embeddings. As evidenced by our dynamic application that requires daily observation of aerosol content in the atmosphere, adding a temporal aspect to location encoders can also be invaluable in applications on dynamic phenomena.
@@
All data uses are openly available at [60] and all software code are available at [61].
@@
Appendix A Evaluation Metrics Tables
Table 2: Within-Region evaluation metrics for five random splits, each with a different training and testing subset. Mean $\pm$ standard deviation are reported. Variation in metrics over the five splits reflects differences in subset difficulty and target distributions.
[…]
Without Lat/Lon $0.73\pm 0.035$ $3.80\pm 0.24$ $0.06\pm 0.72$
With Lat/Lon $0.74\pm 0.016$ $3.79\pm 0.11$ $0.07\pm 0.79$
With Sinusoidal (Lat/Lon) $0.75\pm 0.024$ $3.69\pm 0.17$ $0.11\pm 0.70$
GeoCLIP Location Embeddings (Hadamard) $0.79\pm 0.010$ $3.41\pm 0.08$ $0.03\pm 0.32$
GeoCLIP Location Embeddings (Concat) $0.73\pm 0.036$ $3.81\pm 0.25$ $0.26\pm 0.93$
SatCLIP Location Embeddings (Hadamard) $0.68\pm 0.077$ $4.15\pm 0.45$ $0.39\pm 1.31$
SatCLIP Location Embeddings (Concat) $0.72\pm 0.024$ $3.89\pm 0.16$ $0.17\pm 0.73$
Table 3: Within-Region evaluation metrics for five different spatial splits (with entire time series of test stations held out), each with a distinct training and testing subset. Mean $\pm$ standard deviation are reported. Variation across splits reflects differences in spatial coverage and target distributions.
[…]
Without Lat/Lon $0.62\pm 0.13$ $4.46\pm 0.81$ $0.05\pm 1.63$
With Lat/Lon $0.64\pm 0.067$ $4.37\pm 0.71$ $0.16\pm 1.04$
With Sinusoidal (Lat/Lon) $0.65\pm 0.074$ $4.27\pm 0.64$ $0.36\pm 0.87$
GeoCLIP Location Embeddings $0.67\pm 0.049$ $4.21\pm 0.65$ $0.28\pm 0.47$
Table 4: Out-of-Region (Checkerboard, $\delta{=}8^{\circ}$): metrics on the two disjoint test partitions 1 and 2. The partitions contain different proportions of test stations—48.59% in partition 1 and 51.41% in partition 2—reflecting the inherently uneven geographic distribution of AQS monitoring sites across the contiguous United States.
[…]
Without Lat/Lon 0.59  |  0.53 4.02  |  5.61 -1.00  |  0.27
With Lat/Lon 0.58  |  0.47 4.06  |  5.98 -0.51  |  -1.42
With Sinusoidal (Lat/Lon) 0.58  |  0.47 4.04  |  5.99 0.42  |  -0.65
GeoCLIP Location Embeddings 0.60  |  0.54 3.97  |  5.53 0.19  |  0.19
Table 5: Out-of-Region (Checkerboard, $\delta{=}16^{\circ}$): metrics on the two disjoint test partitions 1 and 2. The partitions contain different proportions of test stations—63.51% in partition 1 and 36.49% in partition 2—reflecting the inherently uneven geographic distribution of AQS monitoring sites across the contiguous United States.
[…]
Without Lat/Lon 0.46  |  0.64 5.62  |  4.01 -1.68  |  -0.15
With Lat/Lon 0.46  |  0.59 5.62  |  4.26 -0.34  |  -1.14
With Sinusoidal (Lat/Lon) 0.37  |  0.58 6.11  |  4.33 -2.19  |  -0.87
GeoCLIP Location Embeddings 0.50  |  0.63 5.40  |  4.09 -0.03  |  0.26